# Towards Generalist Robot Policies: What Matters in Building Vision-Language-Action Models

### v3

@@ Abstract
Foundation Vision Language Models (VLMs) exhibit strong capabilities in multi-modal representation learning, comprehension, and reasoning. By injecting action components into the VLMs, Vision-Language-Action models (VLAs) can be naturally formed and also show promising performance. Existing work has demonstrated the effectiveness and generalization of VLAs in multiple scenarios and tasks. Nevertheless, the transfer from VLMs to VLAs is not trivial since existing VLAs differ in their backbones, action-prediction formulations, data distributions, and training recipes. This leads to a missing piece for a systematic understanding of the design choices of VLAs. In this work, we disclose the key factors that significantly influence the performance of VLA and focus on answering three essential design choices: which backbone to select, how to formulate the VLA architectures, and when to add cross-embodiment data. The obtained results convince us firmly to explain why we prefer VLA and develop a new family of VLAs, RoboVLMs, which require very few manual designs and achieve a new state-of-the-art performance in three simulation tasks and real-world experiments. Through our extensive experiments, which include over 8 VLM backbones, 4 policy architectures, and over 600 distinct designed experiments, we provide a detailed guidebook for the future design of VLAs. In addition to the study, the highly flexible RoboVLMs framework, which supports easy integrations of new VLMs and free combinations of various design choices, is made public to facilitate future research. We open-source all details, including codes, models, datasets, and toolkits, along with detailed training and evaluation recipes at: robovlms.github.io.
Index Terms:
Robot Foundation Models, Vision-Language-Action Models, Generalist Robot Policies
@@ I Introduction
Building generalizable robot policies capable of perceiving, reasoning, and interacting with the physical environment given human instructions has been a long-standing challenge in robotics [5, 7, 4, 35]. Recently, there has been an active exploration into learning robot foundation models by fine-tuning the Vision-Language Models (VLMs) on robot data with certain architectural adjustments. The resulting models, also referred to as Vision-Language-Action Models (VLAs), show promising results in both simulated and real-world tasks [7, 22, 24]111Although the rigorous definition of VLAs is not consistent in different works, we regard fine-tuning pre-trained VLMs as the key factor to identify VLAs in this work.. Except for VLAs, there also exist various generalist policies, e.g., the ones from video models or even from scratch. Therefore, a natural question arises: Why do we prefer VLAs built upon large-scale pre-trained VLMs? Compared with other generalist policies, a mostly believed reason for utilizing VLM-based VLAs is that VLMs have demonstrated strong capabilities in learning generalized and robust representations of multi-modal data, such as text, images/videos, through extensive training on web-scale data. Such capabilities can inspire the adaptation of robot foundation models to bridge the gap between highly diverse open-world scenes and limited robotic data. However, it remains an open problem to what extent large-scale vision-language pre-train facilitates generalist robot policies. Moreover, a large and diverse set of different VLMs emerged rapidly with different kinds of LLM backbone, training data, model sizes, architectures, and training recipes. Which kind of VLM backbones is more suitable for robot manipulation is also a crucial issue for the development of successful VLAs.
@@
Beyond the diversity of different backbones, for generalist robot policies, including VLAs, the structures are more complex and vary in form. Based on the most prevalent existing work [24, 39, 47, 7, 22, 35, 34, 20, 55, 4], we propose a categorization based on 1) how the history and action information are incorporated in VLAs and 2) whether the action space is continuous or discrete. As shown in Fig.2, four types of structure formulations are considered. For history information modeling, two forms are identified: 1) one-step modeling, which utilizes only the current state or observation to produce actions; and 2) history modeling, which processes a sliding window of historical states or observations. Regarding the aggregation of history information, we classify it into two approaches: a) interleaved modeling, which integrates historical observation and action sequences in an interleaved format; and b) policy head, which separately processes each historical step and fuses the information at a distinct policy head for action prediction. Different structures leverage the pre-trained VLMs in different ways. Hence, they may have different features in terms of robustness, generalization ability, and data efficiency when faced with different types of environments and tasks. Therefore, it is practically important but underexplored to understand: How should we formulate VLAs to sufficiently leverage the power of VLMs in practice?
@@
In addition to the VLA itself, the quality and diversity of the training data used to develop VLAs are equally critical. With recent progress achieved by well-known VLAs [39, 7, 22, 35, 4], large-scale data from different sources is important to further improve performance in terms of robustness and generalization against out-of-distribution tasks and environments. Yet, they differ largely in detailed training recipes: some utilize additional data to further pre-train VLMs, refining representations closer to robotic manipulation tasks [4], while others co-train VLAs alongside in-domain tasks [39, 7, 22, 35]. Moreover, by sufficiently pre-trained on diverse manipulation skills, robot policies are expected to learn new skills with minimal demonstrations [13]. Consequently, in the case of developing efficient VLAs, When to leverage the large-scale cross-embodiment data becomes an intriguing issue.
@@
To thoroughly study the aforementioned issues and find the most effective solution for VLAs, our study chose 4 VLA structures, 8 various backbones, and 3 different training data recipes to train the VLA models. In our experiments, we propose a new framework, RoboVLMs, to easily transfer the VLMs into VLAs and implement a fair comparison. We evaluate these models on two popular robot manipulation benchmarks in simulation: CALVIN [32] and SimplerEnv [37]. Moreover, we also trained and evaluated the built VLAs on a self-collected real-world robot manipulation dataset, consisting of 100 manipulation tasks and a total of 74K trajectories. Specifically, we initially selected three commonly used VLMs–LLaVA, Flamingo, and KosMos, as backbones, combining each with the four VLA structures to examine the effects of action space, observation horizon, and history aggregating methods. With the finding that the policy head modeling with continuous action space performs best, we compare 8 various VLMs as the backbone with policy head formulation to answer which backbone is more suitable. Meanwhile, we compare the generalization and data efficiency of different VLA structures. For the question of when to leverage cross-embodiment data, we compare pre-training (the VLAs trained with Open X-Embodiment), finetuning (the VLAs trained with target dataset), and post-training (the VLAs pre-trained with Open X-Embodiment and further finetuned with target dataset). Finally, to confirm the real-world applicability of the VLAs with the optimal configuration, we trained and evaluated them in real-world robot manipulation scenarios, demonstrating generalization across 1) unseen distractors, 2) unseen backgrounds, 3) unseen target objects, and 4) novel skill descriptions.
@@
Through our extensive and comprehensive studies, we derive important insight into building high-performance VLAs around the following questions:

### v4

# What Matters in Building Vision-Language-Action Models for Generalist Robots

@@ Abstract
To utilize Foundation Vision Language Models (VLMs) for robotic tasks and motion planning, the community has proposed different methods for injecting action components into VLMs and building the Vision-Language-Action models (VLAs). In this work, we disclose the key factors that significantly influence the performance of VLA on robot manipulation problems and focus on answering three essential design choices: which backbone to select, how to formulate the VLA architectures, and when to add cross-embodiment data. The obtained results convince us firmly to explain why we prefer VLA and develop a new family of VLAs, RoboVLMs, which require very few manual designs and achieve a new state-of-the-art performance in three simulation tasks and real-world experiments. Through our extensive experiments, which include over 8 VLM backbones, 4 policy architectures, and over 600 distinct designed experiments, we provide a detailed guidebook for the future design of VLAs. In addition to the study, the highly flexible RoboVLMs framework, which supports easy integrations of new VLMs and free combinations of various design choices, is made public to facilitate future research. We open-source all details, including codes, models, datasets, and toolkits, along with detailed training and evaluation recipes at: robovlms.github.io.
@@ I Introduction
Building generalizable robot policies capable of perceiving, reasoning, and interacting with the physical environment given human instructions has been a long-standing challenge in robotics [5, 6, 4, 33, 27]. Recently, there has been an active exploration of Vision-Language-Action Models (VLAs), namely, a family of learning-based robot foundation models directly built upon generalist Vision-Language Models (VLMs), and they have demonstrated promising results of semantic generality in both simulated and real-world tasks [6, 20, 22] (Although the rigorous definition of VLAs is not consistent in different works, we regard the ones directly built upon large-scale VLMs along with robot-domain data as the key factor to identify VLAs in this work.). As shown in Fig. 1a, existing robot policies can be categorized into four main groups: (1) Model-Free Learning, which encode the state into a latent representation and use a policy network to predict actions, enabling generalization across different embodiments; (2) Model-Based Learning, which rely on the robot’s affordances and explicit models of the robot and environment, limiting their applicability to specific configurations; (3) World Model Based Learning, which predicts future goal images and derives actions via inverse dynamics models; and (4) Vision-Language-Action (VLA) Model, representing a specialized branch of model-free learning that leverages the generalization and robustness of Vision-Language Models (VLMs) by employing them as state encoders. Therefore, a natural question arises: Why do we prefer VLAs built upon large-scale pre-trained VLMs? As shown in Fig. 1a, VLA is a newly raised model-free learning framework that receives natural language instruction as input and learns a generalist policy, rather than task-specific policies in traditional model-free learning. The widely-believed reason is that VLAs may inherit the power of VLMs, which have been shown powerful in learning generalized and robust representations of multi-modal data, such as text, images/videos, through extensive training on web-scale data. Such capabilities can facilitate the adaptation of robot foundation models to highly diverse open-world scenes with limited robotic data. However, it remains an open problem how large-scale vision-language pre-training facilitates generalist robot policies.
While VLAs have shown early promise, effectively transferring pre-trained VLMs into high-performing robot policies remains non-trivial. Challenges still remain in how to sufficiently and efficiently select and utilize: 1) vision-language backbones; 2) formulation that best leverages multi-modal representations; and 3) the increasing robot data. As shown in Fig. 2, to address these challenges, we focus on three critical questions:
1. Which kind of VLM backbones are suitable for VLAs? Selecting an appropriate VLM backbone is a fundamental yet underexplored design choice in adapting vision-language models for robotic control. Modern VLMs vary widely in architecture. They differ in model capacity, visual encoder structure, fusion mechanisms, and data scale. While these models are originally trained for general vision-language tasks, their effectiveness in policy learning depends on how well they can encode physical object properties, spatial relationships, and task-relevant semantics. Existing works are based on diversified VLM structures, and most of them show promising results on adaptation to diverse and complicated manipulation skills. Yet, there is no comprehensive study on how backbone choices impact downstream robot manipulation performance, and hence, it remains an open problem: Which kind of VLM backbones facilitates the VLA learning most.
@@
2. How should VLAs be formulated to best leverage VLMs? Beyond the diversity of different backbones, for generalist robot policies, including VLAs, the structures are more complex and vary in form. Based on the most prevalent existing work [22, 38, 45, 6, 20, 33, 32, 18, 53, 4], we propose a categorization based on 1) how the history and action information are incorporated in VLAs and 2) whether the action space is continuous or discrete. As shown in Fig. 1b, four types of structure formulations are considered. For history information modeling, two forms are identified: 1) one-step modeling, which utilizes only the current state or observation to produce actions; and 2) history modeling, which processes a sliding window of historical states or observations. Regarding the aggregation of history information, we classify it into two approaches: a) interleaved modeling, which integrates historical observation and action sequences in an interleaved format; and b) policy head, which separately processes each historical step and fuses the information at a distinct policy head for action prediction. These above-mentioned formulations leverage the pre-trained VLMs in their different ways. Hence, they may have different features in terms of robustness, generalization ability, and data efficiency when faced with various types of environments and tasks. Therefore, it is practically important but underexplored to understand: How should we formulate VLAs to sufficiently leverage the power of VLMs in practice?
@@
3. When should we leverage the public diverse robot data for optimal training? In addition to the VLA itself, the quality and diversity of the training data used to develop VLAs are equally critical. With recent progress achieved by well-known VLAs [38, 6, 20, 33, 4], large-scale data from different embodiments (also known as cross-embodiment data) is important to further improve performance in terms of robustness and generalization against out-of-distribution tasks and environments. Yet, they differ largely in detailed training recipes: some utilize cross-embodiment data to 1) pre-train a foundation VLA from VLMs by refining representations closer to robotic manipulation tasks [4, 54] and 2) post-train the VLA with in-domain data; while others co-train VLAs with cross-embodiment data alongside in-domain data [38, 6, 20, 33]. Moreover, by being sufficiently pre-trained on cross-embodiment datasets, VLAs are expected to learn new skills with minimal demonstrations [12]. Consequently, in the case of developing efficient VLAs, when to leverage the large-scale cross-embodiment data becomes an intriguing issue.
@@
To thoroughly study the aforementioned key questions, we propose a general framework, RoboVLMs, to easily transfer the VLMs into VLAs and implement a fair comparison. We evaluate these models on two popular robot manipulation benchmarks in simulation: CALVIN [30] and SimplerEnv [35]. Moreover, we also trained and evaluated the built VLAs on a real-world robot manipulation dataset, consisting of 100 manipulation tasks and a total of 74K trajectories. Specifically, we initially selected three commonly used VLMs–LLaVA, Flamingo, and KosMos, as backbones, combining each with the four VLA structures illustrated in Fig. 1b to examine the effects of action space, observation horizon, and history aggregating methods. With the finding that the policy head modeling with continuous action space performs best, we compare 8 various VLMs as the backbone with policy head formulation to answer which backbone is more suitable. Meanwhile, we compare the generalization and data efficiency of different VLA structures. For the question of when to leverage cross-embodiment data, we compare co-training (the VLAs trained with Open X-Embodiment), finetuning (the VLAs trained with the target dataset), and post-training (the VLAs co-trained with Open X-Embodiment and further finetuned with the target dataset). Finally, to confirm the real-world applicability of the VLAs with the optimal configuration, we trained and evaluated them in real-world robot manipulation scenarios, demonstrating generalization across 1) unseen distractors, 2) unseen backgrounds, 3) unseen target objects, and 4) novel skill descriptions.
@@
Through our extensive and comprehensive studies, we derive conclusions on how to build high-performance VLAs around the following questions: